OpenStudy (anonymous):

The Jones family paid $150 to a painting contractor to stain their 12-by-15-foot deck. The Smiths have a similar deck that measures 16 ft by 20 ft. What price should the Smith family expect to pay to have their deck stained?

OpenStudy (watchmath):

320/180 * 150

OpenStudy (anonymous):

Find the price per square foot from the Jones' deck and multiply that by the area of the Smith's deck.

OpenStudy (anonymous):

you get 150/180 as the price per square foot, then multiply this by 320 (square footage of Smith deck)

OpenStudy (anonymous):

so what the problem that i have to wite?

OpenStudy (anonymous):

so 150/180=320/x right

OpenStudy (anonymous):

150/180=x/320 price is on top

OpenStudy (anonymous):

i got 266.66

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

so can i round it up?

OpenStudy (anonymous):

if you want to be exact it is 800/3 or 266 2/3

OpenStudy (anonymous):

if you want an estimate you could round to $266.67

OpenStudy (anonymous):

oh ok and can u help me with http://openstudy.com/groups/mathematics/updates/4dd415b2d95c8b0b0de555c4

OpenStudy (anonymous):

sure let me look

OpenStudy (watchmath):

You are a good TA rsvitale :). Deserve a medal :).